OpenStudy (anonymous):

What is the x and y- intercept (s) , for the points (2, -3) and (-8, 2)

OpenStudy (anonymous):

For the points or the line that forms when you connect the points? Because the points themselves are not intercepts, they're simply points.

OpenStudy (anonymous):

i don't know how to

OpenStudy (anonymous):

Well you can find the slope of the line and use that slope and one of the points given to create the equation of this line. With the equation, you can find the intercepts easily. :) Points: (2, -3) = (x1, y1) (-8, 2) = (x2, y2)

OpenStudy (anonymous):

Once you find the slope, use the equation: \[y = m(x-p)+q\] where m is the slope and (p, q) is one of the points, i.e. (2, -3).

OpenStudy (anonymous):

Then to find the intercepts, think about when you have an x-intercept and when you have a y-intercept.  The y-intercept is found when x=0 (no right/left movement). Similarly, the x-intercept is found when y=0 (no up/down movement).

OpenStudy (anonymous):

ok. thank you so much

OpenStudy (anonymous):

Hope it works out.